# Estimating Heterogeneous Treatment Effects by Combining Weak Instruments and Observational Data

### v1

@@ 4.1 Integrating Observational and Experimental Data via Parametric Extrapolation
2:  Learn $\widehat{\tau}^{O}(x)$ using ${\mathcal{T}}$ on $O$. Let $\widetilde{\textbf{Y}}\in\mathbb{R}^{d},\widetilde{\textbf{X}}\in\mathbb{R}^{n\times d}$.
@@
We consider parametric classes $\mathcal{B}_{\phi}=\{\theta^{T}\phi(x):\theta\in\mathbb{R}^{d}\}$ for a known mapping $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$. Since the compliance factor $\gamma(x)$, instrument propensity $\pi_{Z}(x)$, and the parameter of interest $\theta^{T}$ are learned from the same dataset $E$, we propose the following $K$-fold cross-fitted estimation procedure:
@@
where $\widehat{w}^{(k)}(X^{E}_{i}):=\widehat{\gamma}^{(k)}(X^{E}_{i})\widehat{V}^{(k)}_{Z}(X^{E}_{i})$, and the compliance factor $\widehat{\gamma}^{(k)}$ and instrument propensity $\pi_{Z}^{(k)}$, $k\in[K]$ are trained on $E$ excluding the $k^{\text{th}}$ fold containing indices $\mathcal{I}_{k}^{E}$. K-fold cross-fitting is crucial because it ensures that the weights are learned from data distinct from that used in the ERM algorithm. This separation is essential for maintaining desirable theoretical properties as we remain methodologically agnostic to the techniques used for learning $\gamma$ and $\pi_{Z}$.
The compliance factor can be estimated using standard machine learning classification algorithms, either by training separate classifiers for $E\mid Z=1$ and $E\mid Z=0$ or by using one classifier with $Z$ as an additional feature. Similarly, instrument propensity estimation is a straightforward classification task with $Z$ as the target. Given estimates $\widehat{\tau}^{O}$, $\widehat{\gamma}^{(k)}$, and $\pi_{Z}^{(k)}$, the result in is obtained by running an OLS procedure with the target $\widetilde{Y}_{i}^{E}-\widehat{w}^{(k)}(X^{E}_{i})\widehat{\tau}^{O}(X^{E}_{i})$ and the design matrix $\widetilde{\textbf{X}}=W(X^{E})\Phi(X^{E})$. Here, $W(X^{E})={\rm diag}(\widehat{\gamma}(X_{1}^{E})\widehat{V}_{Z}(X_{1}^{E}),\dots,\widehat{\gamma}(X^{E}_{n_{E}})\widehat{V}_{Z}(X^{E}_{n_{E}}))$, and $\Phi(X^{E})=(\phi(X_{1}^{E}),\dots,\phi(X^{E}_{n_{E}}))^{T}$. The two-step procedure is detailed in Algorithm 1.
@@
    (Instrument Positivity) $\epsilon\leq\pi_{Z}(X^{E}),\widehat{\pi}_{Z}(X^{E})\leq 1-\epsilon$ for some $\epsilon>0$.
@@
    (Realizability of $b(x)$) $b(x)\in\mathcal{B}_{\phi}$, i.e. $\tau(x)-\tau^{O}(x)=\theta^{T}\phi(x)$.
@@
Remark 1 (Instrument Propensity).
While we describe our estimation technique in full generality, in practice, the instrument propensity $\pi_{Z}(x)$ is often known, which significantly streamlines both our estimation procedure and our theoretical results. Specifically, if $\pi_{Z}(x)$ is known, there is no need to learn this function from data, and the algorithm’s convergence rates simplify to $O_{p}\left(r_{\gamma}(n_{E})+r_{\tau^{O}}(n_{O})+1/\sqrt{n_{E}}\right)$, which are solely dependent on the rates of the biased CATE $\widehat{\tau}^{O}$ and the compliance term $\widehat{\gamma}$.
We provide the proof of Theorem 2 in Appendix B. The key idea is that weighted OLS remains consistent if the estimates for $\widehat{\gamma}$, $\widehat{\pi}_{Z}$, and $\widehat{\tau}^{O}$ are consistent. However, the convergence rate is limited by the slowest rate among them. Typically, $\pi_{Z}$ is known, so the convergence rate is mainly influenced by the rate of $\widehat{\gamma}$ and $\widehat{\tau}^{O}$. This underscores the trade-off in using both datasets to accurately estimate effects on the target population.
@@ 4.2 Integrating Observational and Experimental Data via a Common Representation
Without expert knowledge, the mapping $\phi(x)$ may not be known a priori. In this section, we describe a method to jointly learn the unbiased CATE and the mapping $\phi(x)$ (referred to as the representation) under the assumption that the true CATE $\tau(x)$ and the biased CATE $\tau^{O}(x)$ share a common representation. This approach leverages machine learning techniques that assume a common structure across tasks, such as multi-task and transfer learning. In causal inference, a shared representation may be assumed between treatment arms [42, 43] or between randomized data and confounded observational data [20]. This framework enables us to learn the bias function $b(x)$ when $\phi(x)\in\Phi$ is unknown.
We consider a class $\Phi$ of representations $\phi(x):\mathcal{X}\rightarrow\mathbb{R}^{d}$ and assume there exists a $\phi\in\Phi$ common to both the true and biased CATEs. Specifically, there are linear hypotheses $h,h^{O}\in\mathbb{R}^{d}$ such that $\tau(x)=h^{T}\phi(x)$ and $\tau^{O}(x)=(h^{O})^{T}\phi(x)$, leading to the bias function $b(x)=(h-h^{O})^{T}\phi(x):=\nu^{T}\phi(x)$. For simplicity, we focus on linear-in-representation classes, but more complex hypotheses $h$ with $\tau(x)=h(\phi(x))$ can be considered – see [42, 20]. Thus, $b(x)\in\mathcal{B}_{\phi}$ for the unknown $\phi$, with $\mathcal{B}_{\phi}$ defined in Section 4.1. Suppose there exists an ERM algorithm ${\mathcal{T}}$ for jointly learning $\phi(x)$ and $h^{O}$ from the observational data, $O$. Our learning algorithm proceeds as follows: first, learn $\widehat{\phi}(x)$ and $\widehat{h}^{O}$ using ${\mathcal{T}}$ with $O$, along with $\widehat{\gamma}^{(k)}(x)$ and $\widehat{V}^{(k)}(x)$ as described in Section 4.1. In the second stage, we use the following ERM procedure to estimate the parameter $\nu$:
We detail this procedure in Algorithm 2. We can then recover $\widehat{\tau}(x)$ by setting $\widehat{\tau}(x)=(\widehat{h}^{O}+\widehat{\nu})^{T}\widehat{\phi}(x)$.
@@
Let $\Phi$ to be a class of feed-forward neural networks. Then $\widehat{\phi}(x),\widehat{h}^{O}$ and $\widehat{\tau}^{O}(x)$ can be jointly learned by composing $\Phi$ with two linear output heads for $Y^{O}\mid A^{O}=1,X^{O}=x$ and $Y^{O}\mid A^{O}=0,X^{O}=x$, respectively. By taking the difference between the two output heads, we can reconstruct $\widehat{\tau}^{O}(x)$, assuming that $\mathbb{E}[Y^{O}\mid A^{O}=1,X^{O}=x]$ and $\mathbb{E}[Y^{O}\mid A^{O}=0,X^{O}=x]$ are also linear in $\phi$ (see [42, 43]). Without this assumption, we can learn $\tau^{O}(x)$ directly by composing $\Phi$ with one linear output layer and considering the pseudo-outcome $\frac{Y^{O}A^{O}}{\pi_{A}(X^{O})}-\frac{Y^{O}(1-A^{O})}{(1-\pi_{A}(X^{O}))}$. Here, $\pi_{A}(X^{O})=P(A^{O}=1\mid X^{O})$ is the treatment propensity in $O$ and can be learned using any black-box machine learning classifier.
@@
We provide the proof of Theorem 3 in Appendix B. This result relies on the realizability assumption in $\Phi$ and the linear-in-representation structure of $\tau$ and $\tau^{O}$. In Example 1, $r_{\phi}(n)$ bounds the generalization error for feed-forward neural networks. For ReLU activations and bounded outputs, $r_{\phi}(n)=C\sqrt{WL\log W\log n}/\sqrt{n}$, where $W$ is the total number of weights, $L$ is the number of layers, and $C$ is a constant independent of $n$ and $W$ [51, 13]. This rate is parametric but scales linearly with $W$, which is problematic for overparameterized networks. For 1-Lipschitz activations and bounded weights, [15] derive a rate of $r_{\phi}(n)=C\sqrt{\Pi_{l=1}^{L}M(l)}/{n^{1/4}}$, where $M(l)$ bounds the Frobenius norm of layer $l$’s weight matrix.
Practical Guidance in High-Dimensional Settings When $\phi(x)$ is high-dimensional, managing the complexity of $\mathcal{B}_{\phi}$ through regularization is critical since we are using the bias function $b(x)$ to extrapolate the CATE beyond low-variance regions where compliance is low and overfitting risks are high. In the parametric extrapolation approach (Section 4.1), implementing $L_{1}$ or $L_{2}$ regularization via Lasso or Ridge regression in the final step effectively controls model complexity. In the shared representation approach in Example 1, regularization not only helps control the parameters $h^{O}$ and $\nu$, but can also prevent over-parametrization in the neural network $\phi$. The choice between $L_{1}$ and $L_{2}$ regularization, and their application, should align with the data-generating process and the desired model characteristics.
@@ 5 Experimental Results
We apply our method to both simulated and real-world data. First, we use the confounded synthetic data example from [26], along with a similar DGP to simulate an IV study, maintaining the same confounding structure and treatment effects. Using this DGP, we evaluate Algorithm 1 and Algorithm 2 in estimating the unbiased CATE by integrating these datasets. Next, we demonstrate our estimators on a real-world dataset examining the impact of 401(k) participation on financial wealth. Details on model implementation, hyperparameter selection, and validation procedures are in Appendix C. The replication code is available at https://github.com/CausalML/Weak-Instruments-Obs-Data-CATE.
@@
Figure 1: Means and standard errors of estimates from 100 simulated dataset pairs $(O,E)$ using Random Forest (top) or Neural Network (bottom) learners. (1(a)): Biased observational CATE $\tau^{O}(x)$. (1(b)): High variance CATEs from the IV dataset using Equation 3. (1(c)): CATEs from Algorithm 2 using parametric extrapolation (top) or representation learning (bottom).
@@ 5.1 Simulation Studies
We generate the observational dataset $O=(X^{O},A^{O},Y^{O})$ as follows (see [26]):
@@
In this DGP, the true CATE is given by $\tau(x)=0.75x^{2}+2x+1$, whereas the biased observational CATE is represented by $\tau^{O}(x)=0.75x^{2}+x+1$. This results in a bias $b(x)=x$, which is linear in $x$. We modify this DGP to generate the experimental IV dataset $E=(X^{E},Z^{E},A^{E},Y^{E})$ as follows:
@@
We generate 100 observational and IV datasets, each with 5000 samples, from the proposed DGP. We apply Algorithm 1 to each dataset. With a randomized instrument, $\pi_{Z}(x)=0.5$. We estimate $\gamma(x)$ as the difference between Random Forest (RF) classifiers trained on $X^{E},A^{E}\mid Z^{E}=0$ and $X^{E},A^{E}\mid Z^{E}=1$. The biased observational CATE is modeled using the T-learner approach [30], with RF regressors trained on $X^{O},Y^{O}\mid A^{E}=0$ and $X^{O},Y^{O}\mid A^{O}=1$. For comparison, we implement a CATE estimator for the experimental data using Equation 3. We compute $\delta_{Y}(x)$ as the difference between RF regressors trained on $X^{E},Y^{E}\mid Z^{E}=0$ and $X^{E},Y^{E}\mid Z^{E}=1$, then divide by $\widehat{\gamma}(x)$, clipping the compliance score at $0.1$. We calculate $\widehat{\gamma}$, $\widehat{\tau}^{O}(x)$, and $\widehat{\tau}^{E}(x)$ for each dataset pair and proceed with the second step of Algorithm 1 by setting $\phi(x)=x$.
@@
In Figure 1 (top row), we show the means and standard errors of our estimators across 100 simulations. The first two plots illustrate the learned observational CATE $\widehat{\tau}^{O}(x)$ and the learned IV CATE $\widehat{\tau}^{E}(x)$. As expected, $\widehat{\tau}^{O}(x)$ shows clear bias, while $\widehat{\tau}^{E}(x)$ has high variance despite aggressive compliance score clipping. The third plot presents results from Algorithm 1, showing that $\widehat{\tau}(x)$ is unbiased and has low variance across $\mathcal{X}$, effectively leveraging the strengths of both datasets to capture the true CATE and address the limitations of each individual study design.
@@
We note that in our DGP, $\tau(x)$, $\tau^{O}(x)$, and $b(x)$ are linear in the polynomial representation $(x,x^{2})$. Thus, we use Algorithm 2 with Example 1 to learn the true CATE and the common representation from the generated dataset. For consistency, we employ feed-forward neural networks (NNs) to estimate all quantities. The estimator for $\widehat{\gamma}$ uses a NN with a sigmoid activation in the output layer, trained on $X^{E}$ with the pseudo-outcome $2A^{E}Z^{E}-2A^{E}(1-Z^{E})$. The representation $\phi(x)$ and the biased CATE $\tau^{O}(x)$ are derived using a representation network with two output heads for learning $Y^{O}\mid X^{O},A^{O}=0$ and $Y^{O}\mid X^{O},A^{O}=1$. A similar dual-head approach is used to learn $\delta_{Y}(x)$, focusing on $Y^{E}\mid X^{E},Z^{E}=0$ and $Y^{E}\mid X^{E},Z^{E}=1$. When calculating $\delta_{Y}(x)/\gamma(x)$, we clip the compliance score at $0.1$. Unlike Algorithm 1, we don’t guarantee the polynomial representation will be fully captured by the chosen representation class, but we expect a sufficiently flexible $\Phi$ to adequately represent these relationships.
@@
The means and standard errors of our estimators from 100 simulations using neural networks and Algorithm 2 are shown in Figure 1 (bottom row). As before, $\widehat{\tau}^{O}(x)$ shows bias, while $\widehat{\tau}^{E}(x)$ has high variance in low-compliance regions, despite score clipping. However, Figure 1(c) shows that $\widehat{\tau}$ remains unbiased with relatively low variance across $\mathcal{X}$. This demonstrates that combining observational and IV data, where the biased and true CATE share a representation, allows us to reliably learn both the representation and the unbiased CATE, overcoming the limitations of each individual study.

### v2

@@ 4.1 Integrating Observational and Experimental Data via Parametric Extrapolation
2:  Learn $\widehat{\tau}^{O}(x)$ using ${\mathcal{T}}$ on $O$. Let $\widetilde{\textbf{Y}}\in\mathbb{R}^{n_{E}},\widetilde{\textbf{X}}\in\mathbb{R}^{n_{E}\times d}$.
@@
We consider a parametric class $\mathcal{B}_{\phi}=\{\theta^{T}\phi(x):\theta\in\mathbb{R}^{d}\}$ for a known mapping $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$. Since the compliance factor $\gamma(x)$, instrument propensity $\pi_{Z}(x)$, and the parameter of interest $\theta^{T}$ are learned from the same dataset $E$, we propose the following $K$-fold cross-fitted estimation procedure:
@@
where $\widehat{w}^{(k)}(X^{E}_{i}):=\widehat{\gamma}^{(k)}(X^{E}_{i})\widehat{V}^{(k)}_{Z}(X^{E}_{i})$, and the compliance factor $\widehat{\gamma}^{(k)}$ and instrument propensity $\widehat{\pi}_{Z}^{(k)}$, $k\in[K]$ are trained on $E$ excluding the $k^{\text{th}}$ fold containing indices $\mathcal{I}_{k}^{E}$. K-fold cross-fitting is crucial because it ensures that the weights are learned from data distinct from that used in the ERM algorithm. This separation is essential for maintaining desirable theoretical properties as we remain methodologically agnostic to the techniques used for learning $\gamma$ and $\pi_{Z}$.
The compliance factor $\gamma(x)=\mathbb{E}[A^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[A^{E}\mid Z^{E}=0,X^{E}=x]$ can be estimated using standard machine learning classification algorithms, either by training separate classifiers for $A^{E}\mid Z^{E}=1,X^{E}=x$ and $A^{E}\mid Z^{E}=0,X^{E}=x$ or by using one classifier with $Z^{E}$ as an additional feature. Similarly, instrument propensity estimation is a straightforward classification task with $Z^{E}$ as the target. Given estimates $\widehat{\tau}^{O}$, $\widehat{\gamma}^{(k)}$, and $\widehat{\pi}_{Z}^{(k)}$, the result in Equation 5 is obtained by performing an OLS procedure with the targets $\widetilde{Y}_{i}^{E}-\widehat{w}^{(k)}(X^{E}_{i})\widehat{\tau}^{O}(X^{E}_{i})$ and the design matrix $\widetilde{\textbf{X}}=W(X^{E})\Phi(X^{E})$. Here, $W(X^{E})={\rm diag}(\widehat{w}^{(k)}(X^{E}_{i}),\dots,\widehat{w}^{(k)}(X^{E}_{n_{E}}))$, and $\Phi(X^{E})=(\phi(X_{1}^{E}),\dots,\phi(X^{E}_{n_{E}}))^{T}$. The two-step procedure is detailed in Algorithm 1.
@@
    (Instrument Positivity in $E$) $\epsilon\leq\pi_{Z}(X^{E}),\widehat{\pi}_{Z}(X^{E})\leq 1-\epsilon$ for some $\epsilon>0$.
@@
    (Realizability of $b(x)$) $b(x)\in\mathcal{B}_{\phi}$, i.e. $\tau(x)-\tau^{O}(x)=\theta^{T}\phi(x)$ for some $\theta\in\mathbb{R}^{d}$.
@@
We include the proof of Theorem 2 in Appendix B. The core insight is that weighted OLS remains consistent as long as the estimates for $\widehat{\gamma}$, $\widehat{\pi}_{Z}$, and $\widehat{\tau}^{O}$ are themselves consistent. However, the overall convergence rate is constrained by the slowest of these rates. In most cases, $\pi_{Z}$ is assumed to be known, meaning the convergence rate is primarily dictated by the rates of $\widehat{\gamma}$ and $\widehat{\tau}^{O}$. This result highlights the trade-off involved in leveraging both datasets to achieve accurate effect estimation for the target population.
Remark 1 (Impact of Realizability Violations).
When realizability does not hold, i.e. $b(x)\notin\mathcal{B}$, our estimator may be inconsistent and exhibit asymptotic bias, proportional to the deviation of the true function from $\mathcal{B}$. Nonetheless, conducting this analysis might still be valuable, as the resulting bias may be smaller than confounding bias in observational estimates or the variance from low compliance in IV studies. Thus, even with uncertain realizability, our method may provide more accurate CATE estimates by effectively balancing bias and variance.
@@ 4.2 Integrating Observational and Experimental Data via a Common Representation
Without expert knowledge, the mapping $\phi(x)$ may not be known a priori. In this section, we introduce a method to jointly learn both the unbiased CATE function and the mapping $\phi(x)$ (hereafter referred to as the representation), based on the assumption that the true CATE $\tau(x)$ and the biased CATE $\tau^{O}(x)$ share a common representation. This approach leverages machine learning techniques that assume a common structure across tasks, such as multi-task and transfer learning. In causal inference, it has been suggested that a shared representation can be assumed between treatment arms [47, 48] or between randomized data and confounded observational data [23]. This framework enables us to learn the bias function $b(x)$ even when the mapping $\phi(x)\in\Phi$ is otherwise unknown.
We consider a class $\Phi$ of representations $\phi(x):\mathcal{X}\rightarrow\mathbb{R}^{d}$ and assume that there exists a shared representation $\phi\in\Phi$ between the true and biased CATEs. Specifically, there exist linear hypotheses $h,h^{O}\in\mathbb{R}^{d}$ such that $\tau(x)=h^{T}\phi(x)$ and $\tau^{O}(x)=(h^{O})^{T}\phi(x)$, resulting in the bias function $b(x)=(h-h^{O})^{T}\phi(x):=\nu^{T}\phi(x)$. For simplicity, we focus on linear-in-representation classes, but more complex hypotheses $h$ with $\tau(x)=h(\phi(x))$ can be considered – see [47, 23]. Thus, $b(x)\in\mathcal{B}_{\phi}$ for the unknown $\phi$, with $\mathcal{B}_{\phi}$ defined in Section 4.1. Suppose there exists an ERM algorithm ${\mathcal{T}}$ that can jointly learn $\phi(x)$ and $h^{O}$ from the observational data, $O$. Our learning algorithm proceeds as follows: first, we use ${\mathcal{T}}$ to learn $\widehat{\phi}(x)$ and $\widehat{h}^{O}$ from $O$, alongside estimates $\widehat{\gamma}^{(k)}(x)$ and $\widehat{V}_{Z}^{(k)}(x)$ from $E$ as described in Section 4.1. In the second stage, we apply the following ERM procedure to estimate the parameter $\nu$:
This procedure is detailed in Algorithm 2. Finally, we recover $\widehat{\tau}(x)$ by setting $\widehat{\tau}(x)=(\widehat{h}^{O}+\widehat{\nu})^{T}\widehat{\phi}(x)$.
@@
Let $\Phi$ be a class of feed-forward neural networks. Then $\widehat{\phi}(x),\widehat{h}^{O}$ and $\widehat{\tau}^{O}(x)$ can be jointly learned by composing $\Phi$ with two linear output heads for $Y^{O}\mid A^{O}=1,X^{O}=x$ and $Y^{O}\mid A^{O}=0,X^{O}=x$, respectively. By taking the difference between the two output heads, we can reconstruct $\widehat{\tau}^{O}(x)$, assuming that $\mathbb{E}[Y^{O}\mid A^{O}=1,X^{O}=x]$ and $\mathbb{E}[Y^{O}\mid A^{O}=0,X^{O}=x]$ are also linear in $\phi$ (see [47, 48]). Without this assumption, we can learn $\tau^{O}(x)$ directly by composing $\Phi$ with one linear output layer and considering the pseudo-outcome $\frac{Y^{O}A^{O}}{\pi_{A}(X^{O})}-\frac{Y^{O}(1-A^{O})}{(1-\pi_{A}(X^{O}))}$. Here, $\pi_{A}(X^{O})=P(A^{O}=1\mid X^{O})$ is the treatment propensity in $O$ and can be learned using any black-box machine learning classifier.
@@
We provide the proof of Theorem 3 in Appendix B. This result hinges on the realizability assumption in $\Phi$ and the linear-in-representation structure of both $\tau$ and $\tau^{O}$. In Example 1, $r_{\phi}(n)$ bounds the generalization error for feed-forward neural networks. For ReLU activations and bounded outputs, $r_{\phi}(n)=C\sqrt{WL\log W\log n}/\sqrt{n}$, where $W$ is the total number of weights, $L$ is the number of layers, and $C$ is a constant independent of $n$ and $W$ [56, 15]. While this rate is parametric, it scales linearly with $W$, which becomes problematic for over-parameterized networks. For 1-Lipschitz activations and bounded weights, [18] derive a rate of $r_{\phi}(n)=C\sqrt{\Pi_{l=1}^{L}M(l)}/{n^{1/4}}$, where $M(l)$ bounds the Frobenius norm of layer $l$’s weight matrix.
Practical Guidance in High-Dimensional Settings: When $\phi(x)$ is high-dimensional, controlling the complexity of $\mathcal{B}_{\phi}$ through regularization is crucial, especially since the bias function $b(x)$ is used to extrapolate the CATE into low-variance regions where compliance is low and the risk of overfitting is high. In the parametric extrapolation approach (Section 4.1), applying $L_{1}$ or $L_{2}$ regularization via Lasso or Ridge regression in the final step is effective for controlling model complexity. In the shared representation approach (Example 1), regularization not only helps control the parameters $h^{O}$ and $\nu$ but also prevents over-parametrization in the neural network $\phi$. The choice between $L_{1}$ and $L_{2}$ regularization, and how they are applied, should be aligned with the data-generating process and the specific characteristics of the model.
@@ 5 Experimental Results
We apply our method to both simulated and real-world data. First, we use the confounded synthetic data example from [29], along with a similar data generating process (DGP) to simulate an IV study, maintaining the same confounding structure and treatment effects. Using this DGP, we evaluate Algorithm 1 and Algorithm 2 in estimating the unbiased CATE by integrating these datasets. Next, we demonstrate our estimators on a real-world dataset examining the impact of 401(k) participation on financial wealth. Additional experiments, as well as details on model implementation, hyperparameter selection, and validation procedures are in Appendix C. The replication code is available at https://github.com/CausalML/Weak-Instruments-Obs-Data-CATE.
@@
Figure 2: Means and standard errors of estimates from 100 simulated dataset pairs $(O,E)$ using Random Forest (top) or Neural Network (bottom) learners. (2(a)): Biased observational CATE $\tau^{O}(x)$. (2(b)): High variance CATEs from the IV dataset using Equation 3. (2(c)): CATEs from Algorithm 2 using parametric extrapolation (top) or representation learning (bottom).
@@ 5.1 Simulation Studies
We generate the observational dataset $O=(X^{O},A^{O},Y^{O})$ as follows (see [29]) [^1] For experimental results using a higher-dimensional version of this DGP, refer to Appendix C.:
@@
In this DGP, the true CATE is given by $\tau(x)=0.75x^{2}+2x+1$, whereas the biased observational CATE is represented by $\tau^{O}(x)=0.75x^{2}+3x+1$. This results in a bias $b(x)=-x$, which is linear in $x$. We modify this DGP to generate the experimental IV dataset $E=(X^{E},Z^{E},A^{E},Y^{E})$ as follows:
@@
We generate 100 observational and IV datasets, each with 5,000 samples, from the proposed DGP. We first apply Algorithm 1 to each dataset. With a randomized instrument, $\pi_{Z}(x)=0.5$. We estimate $\gamma(x)$ as the difference between Random Forest (RF) classifiers trained on $(X^{E},A^{E})\mid Z^{E}=0$ and $(X^{E},A^{E})\mid Z^{E}=1$, i.e. one is trained the subset of data where the instrumental variable $Z^{E}=0$ (using $X^{E}$ and $A^{E}$ as inputs), and another on the subset where $Z^{E}=1$. The biased observational CATE is modeled using the T-learner approach [33], with RF regressors trained on $X^{O},Y^{O}\mid A^{E}=0$ and $X^{O},Y^{O}\mid A^{O}=1$. For comparison, we implement a CATE estimator for the experimental data using Equation 3. We compute $\delta_{Y}(x)$ as the difference between RF regressors trained on $X^{E},Y^{E}\mid Z^{E}=0$ and $X^{E},Y^{E}\mid Z^{E}=1$, then divide by $\widehat{\gamma}(x)$, clipping the compliance score at $0.1$. We calculate $\widehat{\gamma}(x)$, $\widehat{\tau}^{O}(x)$, and $\widehat{\tau}^{E}(x)$ for each dataset pair and proceed with the second step of Algorithm 1 by setting $\phi(x)=x$.
@@
In Figure 2 (top row), we depict the means and standard errors of our estimators across 100 simulations. The first two plots illustrate the learned observational CATE $\widehat{\tau}^{O}(x)$ and the learned IV CATE $\widehat{\tau}^{E}(x)$. As expected, $\widehat{\tau}^{O}(x)$ shows clear bias, while $\widehat{\tau}^{E}(x)$ has high variance despite aggressive compliance score clipping. The third plot presents the results from Algorithm 1, showing that the resulting $\widehat{\tau}(x)$ is both unbiased and has low variance across $\mathcal{X}$. These findings demonstrate that our two-stage estimation procedure effectively leverages the strengths of both datasets to capture the true CATE and address the limitations of each individual study design.
@@
We note that in our DGP, $\tau(x)$, $\tau^{O}(x)$, and $b(x)$ are linear in the polynomial representation $(x,x^{2})$. Thus, we next apply Algorithm 2 with Example 1 to learn the true CATE and the common representation from the generated dataset. For consistency, we employ feed-forward neural networks (NNs) to estimate all quantities. The estimator for $\widehat{\gamma}$ uses a NN with a sigmoid activation in the output layer, trained on $X^{E}$ with the pseudo-outcome $2A^{E}Z^{E}-2A^{E}(1-Z^{E})$. The representation $\phi(x)$ and the biased CATE $\tau^{O}(x)$ are learned using a representation network with two output heads for learning $Y^{O}\mid X^{O},A^{O}=0$ and $Y^{O}\mid X^{O},A^{O}=1$. A similar dual-head approach is used to learn $\delta_{Y}(x)$, by modeling $Y^{E}\mid X^{E},Z^{E}=0$ and $Y^{E}\mid X^{E},Z^{E}=1$ simultaneously. When calculating $\delta_{Y}(x)/\gamma(x)$, we clip the compliance score at $0.1$. Unlike Algorithm 1, we don’t guarantee the polynomial representation will be fully captured by the chosen representation class, but we expect a sufficiently flexible $\Phi$ to adequately represent these relationships.
@@
The means and standard errors of our estimators from 100 simulations using neural networks and Algorithm 2 are shown in Figure 2 (bottom row). As before, $\widehat{\tau}^{O}(x)$ shows bias, while $\widehat{\tau}^{E}(x)$ has high variance in low-compliance regions, despite compliance score clipping. However, Figure 2(c) shows that the $\widehat{\tau}$ returned by Algorithm 2 remains unbiased with relatively low variance across $\mathcal{X}$. This demonstrates that combining observational and IV data, where the biased and true CATE share a representation, allows us to reliably learn both the representation and the unbiased CATE, overcoming the limitations of each individual study.